# Group Distributionally Robust Dataset Distillation with Risk Minimization

### v2

@@ 4.5 Ablation Study
Our focus is primarily on two aspects of utilizing the CVaR loss, i.e. the maximum value and average value of CVaR losses across all sub-clusters. The CVaR loss operates as a complement to the standard cross-entropy optimization. Hence the baseline case, involving only cross entropy loss, mirrors the performance of GLaD. Compared with the maximum value, the average CVaR loss proves more effective in enhancing the validation performance when applied independently. While both average and maximum CVaR loss yield considerable improvement over the baseline, their combined application further fortifies the robustness of the distilled data, which is selected as our eventual implementation.
4.6 Results on More ImageNet Sub-sets
We further provide additional results on ImageNet sub-sets in comparison with GLaD in Tab. 5. The sub-set division remains consistent with that in (Cazenavette et al., 2023). Our proposed robust DD method consistently outperforms the baseline, particularly in terms of the worst accuracy across clustered testing sub-sets. This observation underscores the stability and versatility of the proposed method.
@@ 4.6 Results on More ImageNet Sub-sets
Figure 3: Synthesized sample visualization comparison between GLaD and our proposed method. The samples are initialized with the same original images and updated with different methods.
4.7 Visualization
To explicitly illustrate the impact of CVaR loss on the distillation results, we visualize the sample distribution comparison in Fig. 2. In the feature space, samples optimized by GLaD tend to form small clusters, while the introduction of robust optimization leads to a more evenly distributed distilled dataset. Note that there is no constraint on the feature distribution applied during the distilling process. The proposed robust optimization involves loss calculation on different data partitions, contributing to better coverage over the original data distribution. The more even distribution observed further affirms the effectiveness of our proposed robust distillation method.
Additionally, we compare the synthesized samples of the same ImageNet bonnet class in the pixel space between the baseline GLaD and our proposed method in Fig. 3. The images are initialized with the same original samples for better comparison. Remarkably, the additional CVaR loss introduces more irregular shapes into the image during optimization. These irregular shapes weaken specific features present in each image while introducing common features of the corresponding class, leading to a more even sample distribution in the latent space.
Figure 4: Parameter analysis on the CVaR ratio $\alpha$.
4.8 Parameter Analysis
An analysis is conducted to evaluate the influence of different CVaR ratio $\alpha$ choices in Fig. 4. We vary the $\alpha$ value from 0.2 to 1.0 to explore different ratios of data for calculating CVaR loss. Both the validation performance on the standard testing set and the minimum cluster accuracy reach their peak at $\alpha$=0.8. Including all the samples ($\alpha$=1.0) introduces certain interruptions for the optimization due to large loss values and results in a slight performance drop. On the other hand, considering only a small portion of samples for CVaR loss loses essential information, leading to performance degradation. Notably, even the worst performance obtained at $\alpha$=0.2 is significantly higher than the GLaD baseline, particularly in terms of the minimum cluster accuracy. This observation strongly supports the effectiveness of our proposed robust dataset distillation method.
@@ 5 Conclusion
This paper explored the intricate relationship between DD and its generalization, with a particular focus on performance across uncommon subgroups, especially in regions with low population density. While existing methods prioritize matching convergence properties with the training dataset, we emphasize the importance of considering the broader population distribution. To address this, we introduced an algorithm inspired by distributionally robust optimization, employing clustering and risk minimization to enhance DD. Our theoretical framework, supported by empirical evidence, demonstrates the effectiveness, generalization, and robustness of our approach across diverse subgroups. By prioritizing representativeness and coverage over training error guarantees, our methodology offers a promising avenue for improving the performance of models trained on synthetic datasets in real-world scenarios, paving the way for enhanced applications of dataset distillation in a variety of settings.
Appendix A Appendix
A.1 Theoretical Details
A.1.1 DRO and Generalization
Here we present several results from the literature that indicate that a DRO on an empirical risk minimization problem exhibits generalization guarantees, i.e., approximately minimizes the population (or expected test dataset) loss. Consider (5), repeated here
@@ A.1.1 DRO and Generalization
    Theorem 3 in (Xu & Mannor, 2012) provides a bound for the test error at the DRO solution.
@@
    Theorem 3.1 in  (Zeng & Lam, 2022) presents a probability that $|\mathbb{E}[F(\theta^{*}_{DRO},\mathcal{Q})]-\mathbb{E}[F(\theta^{*}_{opt},\mathcal{Q})]|\leq\epsilon$ as a function of $\epsilon$, where we denote the DRO and the exact minimum, respectively.
A.1.2 Large Deviations and Data Driven DRO
The analysis of the theoretical convergence and robustness properties of our method will rely significantly on the theoretical foundations of data driven DRO established in (Van Parys et al., 2021). To this end, we review a few pertinent definitions. A predictor is a function $c:\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ that defines a model as applied to a data distribution. A data driven predictor uses an empirical distribution of samples ^$\hat{\mathbb{P}}_{T}$ in the prediction.
@@ A.1.2 Large Deviations and Data Driven DRO
In (Van Parys et al., 2021), they prove that the distributionally robust predictor solves (8) in Theorem 4, for discrete distributions, and Theorem 6, for continuous ones.
@@
A.2 Proof of Proposition 3.3
@@
$\begin{array}[]{l}\mathop{CVaR}[F(\theta;c_{i})]=-\frac{1}{\alpha}\left\{\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}F(\theta;z_{t})1[F(\theta;z_{t})\leq f_{\alpha}]\right.\\ \qquad\qquad\left.+f_{\alpha}(\alpha-\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}1[F(\theta;z_{t})\leq f_{\alpha}]\right\}\end{array}$
@@
Step 2: Notice that as long as $\alpha<\frac{1}{2}$, then minimizing the CVaR will also minimize (10).
@@
Appendix B Visualization of Synthesized Samples

### v3

@@ 4 Numerical Results
We conduct an ablation study on the incorporation of CVaR loss in Table 4(b). In addition to CIFAR-10, we also report results on the ImageNet-A sub-set according to the setting in the work of Cazenavette et al. (2023). Both accuracy on the standard testing set and the worst sub-set accuracy “Cluster-min” are presented. Our focus is primarily on two aspects of utilizing the CVaR loss, i.e. the maximum value and average value of CVaR losses across all sub-clusters. The CVaR loss operates as a complement to the standard cross-entropy optimization. Hence the baseline case, involving only cross entropy loss, mirrors the performance of GLaD. Compared with the maximum value, the average CVaR loss proves more effective in enhancing the validation performance when applied independently. While both average and maximum CVaR loss yield considerable improvement over the baseline, their combined application further fortifies the robustness of the distilled data, which is selected as our eventual implementation. Besides, cross entropy loss is still employed in the implementation for a stable optimization.
Table 5: Top-1 robustness evaluation on IDM (CIFAR-10) and GLaD (TinyImageNet). Setting Method IDM IDMR GLaD GLaDR Acc 67.5±0.2 68.1±0.1 24.9±0.5 26.8±0.6 Cluster-min 62.1±0.3 63.2±0.2 20.5±0.8 22.6±0.6 Noise 61.8±0.5 62.8±0.6 22.1±0.6 24.4±0.7 Blur 54.5±0.5 56.2±0.4 20.8±0.3 23.2±0.5 Invert 20.3±0.5 21.5±0.7 10.2±1.2 11.7±0.9 [Uncaptioned image] Figure 2: Analysis on CVaR ratio $\alpha$.
Experiments on Distribution Matching
In addition to methods constrained by gradient matching, the proposed robust DD can also be plugged into DD methods with other matching metrics. We evaluate the efficacy with IDM (Zhao et al., 2023) as the baseline, where distribution similarity is used as the matching metric. Similar to the integration in gradient matching, the CVaR loss is adopted during model training phases. As shown in the first two columns in Table 5, our proposed robust optimization achieves improvement across all metrics. The results demonstrate the possibility of applying RDD to broader dataset distillation methods for robustness enhancement.
Scalability
We also scale up the proposed robust optimization to TinyImageNet, which contains 100 classes, and hence is more challenging compared with other benchmarks. The experiments are conducted on GLaD, and the results are shown in the last two columns of Table 5. When the class number to be optimized is increased, the proposed RDD method provides consistent improvement on the robustness of the distilled data by enhancing model training phases.
Parameter Analysis
An analysis is conducted to evaluate the influence of different CVaR ratio $\alpha$ choices of equation 3 in Figure 2. We vary the $\alpha$ value from 0.2 to 1.0 to explore different ratios of data for calculating CVaR loss. Both the validation performance on the standard testing set and the worst sub-set accuracy (Cluster-min) reach their peak at $\alpha$=0.8. Including all the samples ($\alpha$=1.0) introduces certain interruptions for the optimization due to large loss values and results in a slight performance drop. On the other hand, considering only a small portion of samples for CVaR loss loses essential information, leading to performance degradation. Notably, even the worst performance obtained at $\alpha$=0.2 is significantly higher than the GLaD baseline, particularly in terms of the Cluster-min metric. This observation strongly supports the effectiveness of our proposed robust dataset distillation method.
@@
Figure 4: Synthesized sample visualization comparison between GLaD and our proposed method. The samples are initialized identically.
Visualization
To explicitly illustrate the impact of CVaR loss on the distillation results, we visualize the sample distribution comparison in Figure 4. In the feature space, samples optimized by GLaD tend to form small clusters, while the introduction of robust optimization leads to a more evenly distributed distilled dataset. Note that there is no constraint on the feature distribution applied during the distilling process. The proposed robust optimization involves loss calculation on different data partitions, contributing to better coverage over the original data distribution. The more even distribution observed further affirms the effectiveness of our proposed robust distillation method.
Additionally, we compare the synthesized samples of the same ImageNet bonnet class in the pixel space between the baseline GLaD and our proposed method in Figure 4. The images are initialized with the same original samples for better comparison. Remarkably, the additional CVaR loss introduces more irregular shapes into the image during optimization. These irregular shapes weaken specific features present in each image while introducing common features of the corresponding class, leading to a more even sample distribution in the latent space.
@@ 5 Conclusion
This paper explores the intricate relationship between DD and its generalization, with a particular focus on performance across uncommon subgroups, especially in regions with low population density. To address this, we introduce an algorithm inspired by distributionally robust optimization, employing clustering and risk minimization to enhance DD. Our theoretical framework, supported by empirical evidence, demonstrates the effectiveness, generalization, and robustness of our approach across diverse subgroups. By prioritizing representativeness and coverage over training error guarantees, the method offers a promising avenue for enhancing the models trained on synthetic datasets in real-world scenarios, paving the way for enhanced applications of DD in a variety of settings.
Reproducibility Statement
We have provided detailed instructions to help reproduce the experimental results of this work. In Section. G we provide the statistics of the datasets used for evaluating the proposed method. In Section. H we provide implementation details on the baseline methods and the hyper-parameter settings of our method. The evaluation metric design of the domain-shift setting is also included. Additionally, we have attached the adopted source code in the supplementary material, which will further help understand the proposed method.
Appendix
The appendix is organized as follows: Sec. A provides more detailed theoretical analysis. Sec. C presents the proof for proposition 3.4. Sec. G and Sec. H contain the dataset details and implementation details, respectively. Sec. D offers more experimental results of the proposed robust dataset distillation method. Section I discusses the broader impact; and finally, Section F presents more sample visualization of the robust dataset distillation method.
Appendix A Theoretical Details
A.1 DRO and Generalization
Here we present several results from the literature that indicate that a DRO on an empirical risk minimization problem exhibits generalization guarantees, i.e., approximately minimizes the population (or expected test dataset) loss. Consider equation 9, repeated here
@@ A.1 DRO and Generalization
    Theorem 3 in the work of Xu & Mannor (2012) provides a bound for the test error at the DRO solution.
@@
    Theorem 3.1 in the work of Zeng & Lam (2022) presents a probability that $|\mathbb{E}[F(\theta^{*}_{DRO},\mathcal{Q})]-\mathbb{E}[F(\theta^{*}_{opt},\mathcal{Q})]|\leq\epsilon$ as a function of $\epsilon$, where we denote the DRO and the exact minimum, respectively.
A.2 Large Deviations and Data Driven DRO
The analysis of the theoretical convergence and robustness properties of our method will rely significantly on the theoretical foundations of data driven DRO established in the work of Van Parys et al. (2021). To this end, we review a few pertinent definitions. A predictor is a function $c:\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ that defines a model as applied to a data distribution. A data driven predictor uses an empirical distribution of samples ^$\hat{\mathbb{P}}_{T}$ in the prediction.
@@ A.2 Large Deviations and Data Driven DRO
Van Parys et al. (2021) prove that the distributionally robust predictor solves equation 15 in Theorem 4, for discrete distributions, and Theorem 6, for continuous ones.
@@
Appendix B Generalize to Broader Dataset Distillation Scenarios
Algorithm 2 Zero-order Dataset Distillation
0:   Input: Initial synthetic set $\mathcal{S}$, distilling objective $L(\mathcal{S})$
0:   0:   Execute: 0:       For epoch $E=1,...$ 0:         Compute with $v_{l}\sim\mathcal{N}(0,I)$: $g(\mathcal{S})\approx\nabla L(\mathcal{S})=\frac{1}{M}\sum\limits_{l=1}^{M}\frac{L(\mathcal{S}+\sigma v_{l})-L(\mathcal{S})}{\sigma}v_{l}$ (17) 0:         Set stepsize $s$ (diminishing stepsize rule, or stepwise decay). For instance: $s=0.1/\sqrt{1+E}$ (18) 0:        Set $\mathcal{S}-sg\to\mathcal{S}$
Existing dataset distillation methods adopt differentiable matching metrics to optimize the synthetic data $\mathcal{S}$. However, there can be circumstances where the gradient of the objective cannot be directly acquired. In this case, dataset distillation can still be performed by zero order approximation of the gradient. Starting with some initial $\mathcal{S}$ and proceed as in Algorithm 2, a standard gradient descent approach can be applied, using a zero order approximation of the gradient of $L$ concerning the dataset $\mathcal{S}$, that is $g(\mathcal{S})\approx\nabla L(\mathcal{S})$, computed only using evaluations of $L(\mathcal{S})$.
We can use any method from Berahas et al. (2022) to obtain a gradient estimate of $g\approx\nabla_{\mathcal{S}}L(\mathcal{S})$, here the featured procedure is Gaussian smoothing. The proposed robust dataset distillation method is still able to be applied towards $L(\mathcal{S})$ for more robust optimization.
Appendix C Proof of Proposition 3.4
@@ Appendix B Generalize to Broader Dataset Distillation Scenarios
$\begin{array}[]{l}\texttt{CVaR}[F(\theta;c_{i})]=-\frac{1}{\alpha}\left\{\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}F(\theta;z_{t})1[F(\theta;z_{t})\leq f_{\alpha}]\right.\\ \qquad\qquad\left.+f_{\alpha}(\alpha-\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}1[F(\theta;z_{t})\leq f_{\alpha}]\right\}\end{array}$
@@
Step 2: Notice that as long as $\alpha<\frac{1}{2}$, then minimizing the CVaR will also minimize equation 19.
@@
Appendix D More Experiments and Analysis
Table 6: Top-1 test accuracy on CUB-200 transferred from ImageNet subsets. All the results are reported based on the average of 5 runs. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Transfer Learning GLaD GLaDR
ImageNet-50 $\rightarrow$ CUB-200 49.2±0.9 54.6±0.6
Table 7: Top-1 test accuracy on more challenging test on the syn data trained model. All the results are reported based on the average of 5 runs. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Group GLaD GLaDR
ImageNet-10 $\rightarrow$ ImageNet-A 56.0±1.1 59.6±0.9
ImageNet-10 $\rightarrow$ ImageNet-B 43.0±0.9 47.2±1.0
D.1 More challenging testing scenarios
In Section 4, we demonstrate that the proposed robust dataset distillation method is able to enhance the robustness of the generated data on shifted or truncated testing sets of the same classes. In this section, we evaluate RDD on more challenging testing cases.
Downstream Fine-tuning
The robustness of a neural network can be tested in transfer learning scenarios (Djolonga et al., 2021). Accordingly, we conduct the experiment of transferring the model trained on distilled data to downstream tasks. More concretely, the model is first pre-trained on distilled data from the 50-class subset of ImageNet (Cazenavette et al., 2023). Then it is fine-tuned on CUB-200 Wah et al. (2011). We use the top-1 accuracy on CUB-200 to evaluate the transferability of the distilled data, and the results are shown in Table 6. The results suggest that the data distilled with RDD applied significantly enhances the transferability of the trained model.
One-shot Direct Transfer
A more challenging case would be directly applying the model trained on the distilled data on completely different classes. As in the original model, a classifier is involved to predict the classification probability, directly applying the classifier to different classes would be infeasible. Thus, we adopt a one-shot retrieval-style evaluation approach instead. In this approach, the model is initially trained using the distilled data of 10 classes and subsequently tested on another set of 10 classes. During testing, the corresponding images are passed through the backbone to obtain embedded features, with which a similarity matrix is computed. We use each sample as a query, and check whether the most similar sample among the remaining samples belongs to the same class and report the top-1 accuracy.
The results are presented in Table 7. For these two groups, the model is firstly trained on distilled ImageNet-10 and then tested on ImageNet-A and ImageNet-B, respectively. The subset split of ImageNet-A and ImageNet-B follows the setting in GLaD (Cazenavette et al., 2023). With the robust optimization applied, the distilled data also shows better robustness on unseen classes, demonstrating the efficacy of the proposed method in enhancing the generalizability of the distilled data.
Table 8: Top-1 test accuracy on standard and robustness testing sets. All the results are reported based on the average of 5 runs. The experiment is conducted with DREAM and PDD on the CIFAR-10 IPC-50 setting. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Method Acc Cluster-min Noise Blur Invert
PDD 67.9±0.2 63.9±0.4 58.2±0.7 48.9±1.1 25.7±0.5
PDDR 68.7±0.6 65.1±0.3 59.4±0.9 50.6±0.9 26.7±0.7
DREAM 69.4±0.4 64.7±0.6 58.8±1.2 50.1±0.8 25.7±0.6
DREAMR 69.7±0.5 65.5±0.4 59.8±0.9 50.7±0.7 26.9±0.8
D.2 Application on More DD methods
In addition to IDC (Kim et al., 2022) and GLaD (Cazenavette et al., 2023), we further conduct experiments on CIFAR-10 with DREAM (Liu et al., 2023) and PDD (Chen et al., 2024) as baselines in Table 8. Similar to the implementation for IDC, we use the extra CVaR criterion during the model updating. By the application of robust optimization, all the reported metrics have been improved, especially on the domain-shifted settings. This result demonstrates the generality of the proposed framework across DD methods.
D.3 Efficiency evaluation
As the proposed robust dataset distillation method involves extra CVaR loss calculation during the model updating, the efficiency issue might be a concern. Accordingly, we record the required time to complete the extra robust optimization on IDC. The baseline requires 70s to finish a loop, while the CVaR loss calculation takes up extra 30s. The robust optimization takes up less than 50% of the original calculation time. The extra time is not overwhelming, but brings significant improvements on the robustness of the distilled data.
Table 9: Top-1 test accuracy on different cluster numbers. All the results are reported based on the average of 5 runs. The experiment is conducted with GLaD on CIFAR-10 under the IPC of 50. R indicates the proposed robust dataset distillation method applied. The best results are marked in bold.
Cluster Acc Cluster-min Noise Blur Invert
5 61.6±0.8 55.1±0.9 54.6±0.7 40.3±0.5 17.5±0.9
10 62.5±0.7 56.6±1.0 55.7±0.7 41.1±0.6 18.0±0.9
20 60.9±0.6 54.3±1.2 52.8±0.4 39.8±0.5 17.3±0.7
30 59.9±0.7 53.2±0.9 51.5±0.6 38.9±0.5 17.1±1.0
D.4 Choice of the number of clusters
The CVaR loss calculation involves separating the training set into several sub-sets. Different settings of sub-set division can have influence on the CVaR optimization effects. In this work, the choice of the number of clusters is primarily based on common settings observed in DD benchmarks. Typically, our method is evaluated across different Images-per-class (IPC) settings, such as 1, 10, and 50. For an IPC below 10, we simply use the IPC as the number of clusters, and treat the synthetic samples as the cluster center. However, when dealing with a larger IPC, too many clusters might potentially result in insufficient samples in each cluster for CVaR loss calculation. To address this concern, we conduct a parameter analysis on the number of clusters specifically under an IPC of 50, where the number of clusters varies from 5 to 30. Synthetic samples of the same number are randomly selected to serve as cluster centers, and sub-sets are accordingly separated based on these centers. The experiment is conducted with GLaD on CIFAR-10, and the results are shown in Table 9.
When the number of clusters increases over IPC, as the total training sample number is fixed, the sample number belonging to each sub-set would decrease. And the insufficient samples from each cluster cause sub-optimal CVaR optimization, leading to a performance drop. By empirical observation, we fix the number of clusters to 10 for large-IPC settings.
Table 10: Top-1 test accuracy on different mini-batch sizes. All the results are reported based on the average of 5 runs. The experiment is conducted with GLaD on CIFAR-10 under the IPC of 10. R indicates the proposed robust dataset distillation method applied. The best results are marked in bold.
Mini-batch Size Acc Cluster-min
64 46.9±0.7 40.3±1.1
128 48.9±0.8 45.1±0.8
256 50.2±0.5 46.7±0.9
512 49.3±0.5 45.7±1.2
D.5 Choice of Mini-batch Size
As the group DRO does require abundant samples for precise calculation, we set the sample number in a mini-batch as 256 in our experiments. We further conduct an analysis on the parameter setting, and the results are listed in Table 10 (on CIFAR-10). When there are limited samples in a mini-batch, the performance will be largely influenced. Applying a mini-batch size around 256 brings a mild influence on the validation performance. The slight performance drop when the mini-batch size is set as 512 might be due to less diversity between different mini-batches.
Table 11: Ablation study on the initialization of synthetic samples. All the results are reported based on the average of 5 runs. R indicates the proposed robust dataset distillation method applied. “Cluster” means that the synthetic samples are initialized with clustering centers.
Dataset IPC GLaD GLaDR GLaDR+Cluster
CIFAR-10 1 28.0±0.8 29.2±0.8 29.4±0.9
10 46.7±0.6 50.2±0.5 50.3±0.6
ImageNet-10 1 33.5±0.9 36.4±0.8 36.5±0.8
10 50.9±1.0 55.2±1.1 55.0±1.2
D.6 Choice of Initialization
As the clusters are separated based on the synthetic samples, different initialization of synthetic samples can influence the effects of CVaR loss calculation. In our experiments, we follow the baseline setting, which is random initialization with real samples for both IDC (Kim et al., 2022) and GLaD (Cazenavette et al., 2023). We accordingly conduct an ablation study to evaluate the robustness of RDD on the initialization. In addition to random sampling, we also test initializing the synthetic samples with clustering centers, which leads to a more even distribution. The results are shown in Table 11, where clustering initialization yields similar results to random initialization. Although random initialization cannot guarantee an even distribution at the beginning, during the subsequent optimization process, the algorithm is still robust enough to handle different initializations and provide stable performance improvement over the baseline.
Table 12: Evaluation of the proposed robust dataset distillation on other domain generalization methods with GLaD. All the results are reported based on the average of 5 runs. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Method CIFAR-10 ImageNet-10
GLaD GLaDR GLaD GLaDR
Baseline 46.7±0.6 50.2±0.5 50.9±1.0 55.2±1.1
MMD 47.1±0.6 51.3±0.5 51.8±1.3 56.5±1.2
RSC 47.9±0.5 52.5±0.5 52.4±1.1 56.8±1.0
HYPO 49.0±0.5 53.2±0.6 53.6±1.2 58.1±1.1
D.7 Evaluation on Domain Generalization Baselines
The proposed robust dataset distillation method mainly focuses on enhancing the robustness of distilled datasets. The method can also be combined with other domain generalization methods to further enhance the robustness of model training. We have included the three generalization methods MMD (Li et al., 2018), RSC (Huang et al., 2020), and HYPO (Bai et al., 2024) in Table 12 to serve as training pipelines. The results suggest that on the one hand, RDD consistently provides performance improvement on different training pipelines. On the other hand, the combination of RDD and other domain generalization methods can further improve the results over the baseline.
Table 13: Validation accuracy on more ImageNet sub-sets in comparison with GLaD. The data is distilled with ConvNet-5 and evaluated with ResNet-10 under the IPC of 10. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Dataset Acc Cluster-min
GLaD GLaDR GLaD GLaDR
ImageNet-A 53.9±0.7 57.5±1.2 40.5±0.9 45.8±0.8
ImageNet-B 50.3±0.9 53.8±0.8 42.9±1.1 47.0±1.0
ImageNet-C 49.2±0.8 51.3±0.6 28.2±0.7 32.8±0.6
ImageNet-D 39.1±0.6 40.9±0.7 27.0±0.8 31.3±0.9
ImageNet-E 38.9±0.8 41.1±0.9 25.8±0.6 30.0±0.7
D.8 Results on More ImageNet Sub-sets
GLaD (Cazenavette et al., 2023) designs experiments on multiple ImageNet sub-sets. We further provide results on these sub-sets in comparison with GLaD in Table 13. The sub-set division remains consistent with that in the work of Cazenavette et al. (2023). Our proposed robust DD method consistently outperforms the baseline, particularly in terms of the worst accuracy across clustered testing sub-sets. This observation underscores the stability and versatility of the proposed method.
Appendix E Related Work
Dataset Distillation
Dataset distillation (DD) seeks to distill the richness of extensive datasets into compact sets of synthetic images that closely mimic training performance (Sachdeva & McAuley, 2023). These condensed images prove invaluable for various tasks, including continual learning (Gu et al., 2023b), federated learning (Liu et al., 2022a; Jia et al., 2023), neural architecture search (Such et al., 2020; Medvedev & D’yakonov, 2021), and semi-supervised learning (Vahidian et al., 2020; Joneidi et al., 2020). Existing DD methodologies can be broadly categorized into bi-level optimization and training metric matching approaches. Bi-level optimization integrates meta-learning into the surrogate image update process with the validation performance as a direct optimizatino target (Zhou et al., 2022; Loo et al., 2023). Conversely, metric matching techniques refine synthetic images by aligning with training gradients (Kim et al., 2022; Liu et al., 2023), feature distribution (Sajedi et al., 2023; Zhao et al., 2023), or training trajectories (Wu et al., 2023; Du et al., 2023) compared to the original images. Data parametrization (Kim et al., 2022; Liu et al., 2022b; Wei et al., 2024) and generative prior (Cazenavette et al., 2023; Gu et al., 2023a; Wang et al., 2023a) are also considered for more efficient DD method construction.
Robustness in Dataset Distillation
To the best of our knowledge, this is the first work on subgroup accuracy specifically, or even DRO generally, for Dataset Distillation. We consider a few classes of related works. In the context of DD, adversarial robustness is another popular notion of robustness, which is simply minimization with respect to the worst possible sample in the support of some perturbation on the data sample, rather than the distribution. Adversarial Robustness is more conservative than DRO, however, in some security-sensitive circumstances, this is a more appropriate framework (Wu et al., 2022; Xue et al., 2024). In the field of knowledge distillation, wherein a more parsimonious model, rather than dataset, is of interest, group DRO has been considered in the work of Vilouras et al. (2023) and Wang et al. (2023b).
The concept of group DRO is considered and an implementation thereof is demonstrated in the work of Sagawa et al. (2019), complemented with a regularization strategy that appears to assist in performance for small population subgroups. The coverage of disparate data distributions is also a concern in Non-IID federated learning. Jiao et al. (2023) present convergence theory for DRO in a federated setting.
Appendix F Visualization of Synthesized Samples